the income elasticity of demand of commodity M, is 0.84 a.interprete the magnitude in the elasticity value b.write the formulae for calculationg such elasticity (define all symbols). c.by how much will demand of commodity M changeif income should decrease by 12.5 percent?


1
Expert's answer
2021-05-31T09:10:19-0400

Solution:

a.). Income elasticity of demand is an economic measure of how responsive the quantity demand for a good or service is to a change in income.


Income elasticity of commodity M of 0.84 means that the commodity is income inelastic and a normal good, since the value is positive. Normal goods whose value is between zero and one are normally referred to as basic or necessity goods, therefore commodity M is a basic good. It is, therefore, a commodity that consumers will purchase regardless of changes in their income levels.


b.). The formula for calculating the income elasticity of demand is as follows:


Income elasticity of demand (YED) ="\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; income}"



Income elasticity of demand (YED) = "=\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 }"


Where: Q2 = New quantity demanded

             Q1 = Old quantity demanded

             I2 = New Income

             I1 = Old Income

           YED = Income elasticity of demand.

 

c.). Income elasticity of demand (YED) = "\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; income}"

 

Income elasticity of demand (YED = 0.84

% Change in income = -12.5%


0.84 = "\frac{\%\;change\; in\; quantity\; demanded}{-12.5\%\ }"


% Change in quantity demanded = "-12.5\% \times 0.84 = -10.5\%"

% Change in quantity demanded = -10.5%

 

The demand for commodity M will decrease by 10.5% when income decrease by 12.5%